OpenStudy (anonymous):

How to integrate (1+1/ln(x))/x?

OpenStudy (anonymous):

I use u-sub where u=ln(x), can this work?

OpenStudy (anonymous):

@zepdrix @dan815 @Compassionate @SolomonZelman

zepdrix (zepdrix):

\[\Large\rm \int\limits \frac{1}{x \ln x}dx\]Ahh the brackets are confusing, is this what we're working with?

zepdrix (zepdrix):

\[\Large\rm \frac{1+\frac{1}{\ln x}}{x}dx\]

OpenStudy (anonymous):

Yes, the second one.

zepdrix (zepdrix):

u=ln x? Hmmm that might work.. let's see.\[\Large\rm \int\limits\frac{1+\frac{1}{\ln x}}{x}dx=\int\limits\left(1+\frac{1}{\ln x}\right)\left(\frac{1}{x}dx\right)\] \(\Large\rm du=\dfrac{1}{x}dx\), yes?

OpenStudy (anonymous):

continue

zepdrix (zepdrix):

Then you plug stuff in:\[\Large\rm \int\limits\left(1+\frac{1}{u}\right)\left(du\right)\]And integrate. Looks pretty straight forward from here :)

OpenStudy (anonymous):

I see, thank you.